# Learning Structured Representations by Embedding Class Hierarchy with Fast Optimal Transport

### v1

@@ Abstract
To embed structured knowledge within labels into feature representations, prior work (Zeng et al., 2022) proposed to use the Cophenetic Correlation Coefficient (CPCC) as a regularizer during supervised learning. This regularizer calculates pairwise Euclidean distances of class means and aligns them with the corresponding shortest path distances derived from the label hierarchy tree. However, class means may not be good representatives of the class conditional distributions, especially when they are multi-mode in nature. To address this limitation, under the CPCC framework, we propose to use the Earth Mover’s Distance (EMD) to measure the pairwise distances among classes in the feature space. We show that our exact EMD method generalizes previous work, and recovers the existing algorithm when class-conditional distributions are Gaussian. To further improve the computational efficiency of our method, we introduce the Optimal Transport-CPCC family by exploring four EMD approximation variants. Our most efficient OT-CPCC variant, the proposed Fast FlowTree algorithm, runs in linear time in the size of the dataset, while maintaining competitive performance across datasets and tasks.
@@ 1 Introduction
Supervised classification problems have been a cornerstone in machine learning. While traditional classification methods treat class labels as independent entities, there has been a growing interest in leveraging semantic similarity among classes to improve model performance while following existing label dependency. This is particularly important in fine-grained classification tasks where labels are inherently living on a hierarchy, such as in the CIFAR (Krizhevsky and Hinton, 2009) and ImageNet (Deng et al., 2009) datasets. To capture the hierarchical relationship between classes, Zeng et al. (2022) made a significant contribution by embedding the tree metric of a given class hierarchy into the feature space. They introduced a regularizer, the Cophenetic Correlation Coefficient (CPCC), to correlate the tree metric with the Euclidean distance in the class-conditioned representations. This approach of encoding hierarchical relationship in the latent space not only led to more interpretable features compared to other hierarchical classification methods (Shen et al., 2023), but also improved some generalization performance across multiple datasets.
Despite the value of this framework for embedding class hierarchies, it presents a notable limitation. Specifically, Zeng et al. (2022) uses the $\ell_{2}$ norm between two class Euclidean centroids as the distance of two class distributions. However, this approach overlooks the difference of samples within the same class, and thus cannot fully capture the global structural information given pairs of class-conditional distributions, especially when they are multi-modal. On the other hand, the use of Optimal Transport (OT) as an alternative to Euclidean centroids has first gained attention in the context of clustering, where Euclidean centroids neglect the intrinsic geometry of data distributions (Cuturi and Doucet, 2014). For the same motivation, we propose addressing the limitation of $\ell_{2}$-CPCC (Zeng et al., 2022), by integrating the OT distance, also known as the Wasserstein distance, and its variants, as a more effective metric for measuring the distance between class distributions in the problem of embedding hierarchical representations.
@@ 2 Preliminaries
The hyperparameter $\lambda$ is the regularization factor which controls the strength of ${\mathcal{R}}$. The goal of the problem is to embed the class hierarchical relationship in the feature space. At a high level, the distance of class-conditioned representations should approximate the tree metric of classes in ${\mathcal{T}}$. This extra constraint can be quantitatively characterized by CPCC as the regularizer term ${\mathcal{R}}$ that we will introduce next.
@@ 2.1 Cophenetic Correlation Coefficient
The Cophenetic Correlation Coefficient (CPCC) (Sokal and Rohlf, 1962) was introduced as a regularization term in Zeng et al. (2022) to embed the class hierarchy ${\mathcal{T}}$ in the feature space. CPCC has the same formulation as Pearson’s correlation coefficient that measures the linear dependence between two metrics $t$ and $\rho$:
@@ 2.2 Exact Earth Mover’s Distance
$\min\limits_{{\bm{P}}}\ \langle{\bm{P}},{\bm{D}}\rangle,\quad\mathrm{\,s.t.\ }{\bm{P}}{\bm{1}}_{m}=a,~{}{\bm{P}}^{\top}{\bm{1}}_{n}=b.$ (3)
@@
When the dimension of the feature is 1, fast computation of the EMD distance exists (Houry et al., 2024). Instead of solving a linear program in Eq.3, when $m=n$, $a_{i}=b_{i}=1/n$ for all $i$, the closed form solution for 1d EMD is: $\text{EMD}_{\text{1d}}(\mu,\nu)=\frac{1}{n}\sum_{i=1}^{n}\|z_{(i)}-z^{\prime}_{(i)}\|$, and $z_{(i)}$ is the $i$-th order statistic of the sequence $\{z_{1},\cdots z_{n}\}$ after sorting. When $m\neq n$, $\text{EMD}_{\text{1d}}$ generalizes to solving inverse CDF of $\mu,\nu$ (Peyré and Cuturi, 2019) which requires approximation. To avoid this approximation, after sorting, we can compute the flow matrix ${\bm{P}}$ (Hoffman, 1963) efficiently via a greedy algorithm in linear time (Alg.1) and use the inner product objective of EMD in Eq.3 to compute the result. This greedy flow matching algorithm that generates the optimal transport plan can be used for any $m,n$, including the special case when $m=n$, and fit any non-uniform weights of $a,b$.
@@ 3.1 OT-CPCC Family
Despite its simplicity, the use of class centroids $\ell_{2}$ distance as a measure of class distances has two limitations. First, since there is no guarantee of the distribution of class conditional features, the class mean difference may not accurately reflect the average pairwise Euclidean distance between all the pairs of sample. For example, in Fig.1, class $v$ is a non-unimodal (bi-modal in this example) distribution, which is very likely to happen when class $v$ is an abstract concept, such as coarse classes like household furniture in CIFAR100 (Krizhevsky and Hinton, 2009) or the collection of diverse dog breeds in ImageNet Deng et al. (2009). The blue $\ell_{2}$ line can be a biased estimation of the average of red lines. We will see later (c.f. Sec.3.3) that the direction of blue vector plays an important role in the feature geometry in $\ell_{2}$-CPCC, and therefore we misrepresent the class distribution in $\ell_{2}$-CPCC. Second, it is known (Linial et al., 1994) that, for certain trees such as a basic star graph with three edges, it is impossible to embed tree metric $t$ into $\ell_{2}$ exactly, no matter how large feature dimension $d$ is. It implies if we use the Euclidean centroids to represent a class, we can never reach the optimal CPCC value $1$ during training, which always cause some loss of hierarchical information in the learning of structured representation.
@@
Proposition 3.1 (EMD reduces to $\ell_{2}$ for Gaussians with same covariance).
@@
Although EMD has many appealing properties, the computational cost of EMD is high, where the time complexity of linear programming is cubic time with respect to the data size. Many following work provides approximation methods to reduce the computational cost, such as Sinkhorn (Cuturi, 2013), Sliced Wasserstein Distance (SWD) (Bonneel et al., 2015), Tree Wasserstein Distance (TWD) (Yamada et al., 2022; Le et al., 2019; Takezawa et al., 2021), and FlowTree (FT) (Backurs et al., 2020) (detailed discussion of approximation methods in App.A). Similarly, we can replace $\rho$ with all other EMD approximation methods, and we call all of them as the OT-CPCC family. As a side remark, since all approximation methods provide a ${\bm{P}}$ satisfying constraints in Eq.3 and since EMD is lower bounded by $\ell_{2}$ (see proof of Prop.3.1 in App.B), all OT-CPCC methods are still lower bounded by $\ell_{2}$.
@@ 3.2 Fast FlowTree
(Backurs et al., 2020) is a tree-based method where the structure of the tree ${\mathcal{T}}$ is learned using QuadTree (Samet, 1984) to approximate the Euclidean space. Once the tree has been built, it firsts compute the optimal flow matrix ${\bm{P}}_{t}$ using $t$ as the ground metric with a linear time bottom-up as shown in Alg.2 (Kalantari and Kalantari, 1995) (see App.Fig.8 in for visualization). Then, it estimates EMD with $\langle{\bm{P}}_{t},{\bm{D}}_{\text{$\ell_{2}$}}\rangle$. Given the tree, the time complexity of computing the FT is $O((m+n)d\log(d\Phi))$ where $\Phi$ is the maximum Euclidean distance of any two feature vectors.
@@
The main computational bottleneck of FT is the construction of the tree. However, in our case, we have access to ${\mathcal{T}}$ given as prior knowledge, and thus we can skip the tree construction step. To apply tree-based EMD approximation methods, instead of building a QuadTree in FT, we first construct an augmented tree by extending the leaves of the original label tree ${\mathcal{T}}$ with samples. Specifically, since each leaf node in ${\mathcal{T}}$ corresponds to a class, we extend each leaf node by adding all samples from the corresponding class, as demonstrated in Figure 2. We set all extended edges to have edge weight $1$, and assign the instance weight $a_{i}$ for each data point. The resulting algorithm that applies Alg.2 to the augmented label tree is called Fast FlowTree (FastFT).
One key insight that allows FastFT to be more efficient than FT is that the tree metric between any two data samples from different classes is always identical in the augmented tree ${\mathcal{T}}$, which allows us to reduce the computation of ${\bm{P}}_{t}$ to 1d OT greedy flow matching using Alg.1. The time complexity of FastFT is $O((m+n)d)$, which is optimal among all OT variants.
@@
Please refer to App.C for proof details of Theorem 3.2. Note that all OT-CPCC methods inherently need $\Omega((m+n)d)$ because we access each of $m+n$ data point once and look at each entry in $d$-dimensions. Additionally, FastFT can be applied to any label tree with no restriction on structure or edge weights. To control the importance of data points, the flexibility of FastFT can be achieved by tuning $a_{i}$.
@@
$\ell_{2}$ $O(d)$checkmark N/A
EMD $O(n(d+n^{2}\log n))$checkmark
Sinkhorn $O(n^{2}dI)$checkmark
TWD $O(n^{2}dI)$checkmark
SWD $O(np(d+\log n))$
FT $O(nd\log(d\Phi))$checkmark
FastFT $\Theta(nd)$checkmark
Table 1: Time complexity comparison of different CPCC methods. Let $b$ be the batch size, $I$ be the number of iterations for iterative methods, $p$ be the number of projections for SWD, $\Phi$ be the maximum Euclidean distance of any two feature vectors. We ignore the factor of $\min(b^{2},k^{2})$ shared for all methods due to batch computation. For simplicity we assume $m=n$. In the batch learning setting, $d$ can be much larger than $n$ particularly for the fine-grained classification problems.
@@
To further understand the behavior of different OT approximation methods, we present experiments on synthetic datasets in Fig.3 (see App.H for more details). We first measured the computation time for different OT methods across datasets of varying sizes. Second, we evaluated the error of different OT approximation methods, along with the $\ell_{2}$ distance of the mean of two datasets, in approximating EMD. Two different scenarios were considered, either both distributions were Gaussian or non Gaussian. Under Gaussian distribution, the OT methods’ results were closely aligned with the $\ell_{2}$ method, except for SWD, which supports Prop.3.1 and implies the similarity in performance between OT and $\ell_{2}$-CPCC under Gaussian distributions. In non-Gaussian distribution scenario, notable differences emerged between OT methods and $\ell_{2}$, where SWD is closer to $\ell_{2}$ measurement. FastFT demonstrated exceptional advantage as an approximation method, outperforming Sinkhorn in both computational speed and precision.
@@ 3.3 Differentiability Analysis of OT-CPCC
Denote ${\bm{P}}^{*}$ as the optimal flow of EMD problem, then $\frac{\partial\rho_{\text{EMD}}}{\partial\theta}={\bm{P}}^{*}\cdot\frac{\partial D}{\partial\theta}$.
@@
The application of Danskin’s relies on the optimality of ${\bm{P}}^{*}$. How about the other approximation methods? The differentiability is not a problem in Sinkhorn as the iterative algorithm only involves matrix multiplication. But for the SWD, TWD, FT, the approximate solutions (the flow matrix ${\bm{P}}$) for these methods are not explicitly formulated as solutions to an optimization problem. Following EMD, except for the Sinkhorn variant, our method is to treat flow matrix ${\bm{P}}$ always as a constant for these approximation methods, and stop the gradient through ${\bm{P}}$. For example, $\frac{\partial\rho_{\text{FT}}}{\partial\theta}\vcentcolon={\bm{P}}_{\text{FT}}\cdot\frac{\partial D}{\partial\theta}$.
To provide a more fine-grained understanding of the proposed method, after showing and defining the differentiability of OT-CPCC, in what follows we analyze the gradients of the proposed CPCC variants and compare them with the $\ell_{2}$-CPCC.

### v2

@@ Abstract
To embed structured knowledge within labels into feature representations, prior work (Zeng et al., 2022) proposed to use the Cophenetic Correlation Coefficient (CPCC) as a regularizer during supervised learning. This regularizer calculates pairwise Euclidean distances of class means and aligns them with the corresponding shortest path distances derived from the label hierarchy tree. However, class means may not be good representatives of the class conditional distributions, especially when they are multi-mode in nature. To address this limitation, under the CPCC framework, we propose to use the Earth Mover’s Distance (EMD) to measure the pairwise distances among classes in the feature space. We show that our exact EMD method generalizes previous work, and recovers the existing algorithm when class-conditional distributions are Gaussian. To further improve the computational efficiency of our method, we introduce the Optimal Transport-CPCC family by exploring four EMD approximation variants. Our most efficient OT-CPCC variant, the proposed Fast FlowTree algorithm, runs in linear time in the size of the dataset, while maintaining competitive performance across datasets and tasks. The code is available at https://github.com/uiuctml/OTCPCC.
@@ 1 Introduction
Supervised classification problems have been a cornerstone in machine learning. While traditional classification methods treat class labels as independent entities, there has been a growing interest in leveraging semantic similarity among classes to improve model performance while following existing label dependency. This is particularly important in fine-grained classification tasks where labels are inherently living on a hierarchy, such as in the CIFAR (Krizhevsky & Hinton, 2009) and ImageNet (Deng et al., 2009) datasets. To capture the hierarchical relationship between classes, Zeng et al. (2022) made a significant contribution by embedding the tree metric of a given class hierarchy into the feature space. They introduced a regularizer, the Cophenetic Correlation Coefficient (CPCC), to correlate the tree metric with the Euclidean distance in the class-conditioned representations. This approach of encoding hierarchical relationship in the latent space not only led to more interpretable features compared to other hierarchical classification methods (Shen et al., 2023), but also improved some generalization performance across multiple datasets.
Despite the value of this framework for embedding class hierarchies, it presents a notable limitation. Specifically, Zeng et al. (2022) uses the $\ell_{2}$ norm between two class Euclidean centroids, or equivalently, maximum mean discrepancy with the linear kernel, as the distance of two class distributions. However, this approach overlooks the difference of samples within the same class, and thus cannot fully capture the global structural information given pairs of class-conditional distributions, especially when they are multi-modal. On the other hand, the use of Optimal Transport (OT) as an alternative to Euclidean centroids has first gained attention in the context of clustering, where Euclidean centroids neglect the intrinsic geometry of data distributions (Cuturi & Doucet, 2014). For the same motivation, we propose addressing the limitation of $\ell_{2}$-CPCC (Zeng et al., 2022), by integrating the OT distance, also known as the Wasserstein distance, and its variants, as a more effective metric for measuring the distance between class distributions in the problem of embedding hierarchical representations.
@@ 2 Preliminaries
The hyperparameter $\lambda$ is the regularization factor which controls the strength of ${\mathcal{R}}$. The goal of the problem is to embed the class hierarchical relationship in the feature space. At a high level, the distance of class-conditioned representations should approximate the tree metric of classes in ${\mathcal{T}}$. This extra constraint can be quantitatively characterized by CPCC as the regularizer term ${\mathcal{R}}$.
@@ 2.1 Cophenetic Correlation Coefficient
The Cophenetic Correlation Coefficient (CPCC) (Sokal & Rohlf, 1962) was introduced as a regularization term in Zeng et al. (2022) to embed the class hierarchy ${\mathcal{T}}$ in the feature space. CPCC has the same formulation as Pearson’s correlation coefficient that measures the linear dependence between two metrics $t$ and $\rho$:
@@ 2.2 Exact Earth Mover’s Distance
$\min\limits_{{\bm{P}}}\ \langle{\bm{P}},{\bm{D}}\rangle,\quad\mathrm{\,s.t.\ }{\bm{P}}{\bm{1}}_{n}=a,~{}{\bm{P}}^{\top}{\bm{1}}_{m}=b.$ (3)
@@
When the dimension of the feature is 1, fast computation of the EMD distance exists (Houry et al., 2024). Instead of solving a linear program in Eq.3, when $m=n$, $a_{i}=b_{i}=1/n$ for all $i$, the closed form solution for 1d EMD is: $\text{EMD}_{\text{1d}}(\mu,\nu)=\frac{1}{n}\sum_{i=1}^{n}\|z_{(i)}-z^{\prime}_{(i)}\|$, and $z_{(i)}$ is the $i$-th order statistic of the sequence $\{z_{1},\cdots z_{n}\}$ after sorting. When $m\neq n$, $\text{EMD}_{\text{1d}}$ generalizes to solving inverse CDF of $\mu,\nu$ (Peyré & Cuturi, 2019) which requires approximation. To avoid this approximation, after sorting, we can compute the flow matrix ${\bm{P}}$ (Hoffman, 1963) efficiently via a greedy algorithm in linear time (Alg.1) and use the inner product objective of EMD in Eq.3 to compute the result. This greedy flow matching algorithm that generates the optimal transport plan can be used for any $m,n$, including the special case when $m=n$, and fit any non-uniform weights of $a,b$.
@@ 3.1 OT-CPCC Family
Despite its simplicity, the use of class centroids $\ell_{2}$ distance as a measure of class distances has two limitations. First, since there is no guarantee of the distribution of class conditional features, the class mean difference may not accurately reflect the average pairwise Euclidean distance between all the pairs of sample. For example, in Fig.1, class $v$ is a non-unimodal (bi-modal in this example) distribution, which is very likely to happen when class $v$ is an abstract concept, such as coarse classes like household furniture in CIFAR100 (Krizhevsky & Hinton, 2009) or the collection of diverse dog breeds in ImageNet Deng et al. (2009). The blue $\ell_{2}$ line can be a biased estimation of the average of red lines. We will see later (c.f. Sec.3.3) that the direction of blue vector plays an important role in the feature geometry in $\ell_{2}$-CPCC, and therefore we misrepresent the class distribution in $\ell_{2}$-CPCC. Second, it is known (Linial et al., 1994) that, for certain trees such as a basic star graph with three edges, it is impossible to embed tree metric $t$ into $\ell_{2}$ exactly, no matter how large feature dimension $d$ is. It implies if we use the Euclidean centroids to represent a class, we can never reach the optimal CPCC value $1$ during training, which always cause some loss of hierarchical information in the learning of structured representation.
@@
Proposition 3.1 (EMD reduces to $\ell_{2}$ between means of Gaussian with same covariance).
@@
Although EMD has many appealing properties, the computational cost of EMD is high, where the time complexity of linear programming is cubic time with respect to the data size. Many following work provides approximation methods to reduce the computational cost, such as Sinkhorn (Cuturi, 2013), Sliced Wasserstein Distance (SWD) (Bonneel et al., 2015), Tree Wasserstein Distance (TWD) (Yamada et al., 2022; Le et al., 2019; Takezawa et al., 2021), and FlowTree (FT) (Backurs et al., 2020) (detailed discussion of approximation methods in App.A). Similarly, we can replace $\rho$ with all other EMD approximation methods, and we call all of them as the OT-CPCC family. As a side remark, since all approximation methods provide a ${\bm{P}}$ satisfying constraints in Eq.3 and since EMD is lower bounded by $\ell_{2}$ (see proof of Prop.3.1 in App.C), all OT-CPCC methods are still lower bounded by $\ell_{2}$.
@@ 3.2 Fast FlowTree
(Backurs et al., 2020) is a tree-based method where the structure of the tree ${\mathcal{T}}$ is learned using QuadTree (Samet, 1984) to approximate the Euclidean space. Once the tree has been built, it firsts compute the optimal flow matrix ${\bm{P}}_{t}$ using $t$ as the ground metric with a linear time bottom-up as shown in Alg.2 (Kalantari & Kalantari, 1995) (see App.Fig.8 in for visualization). Then, it estimates EMD with $\langle{\bm{P}}_{t},{\bm{D}}_{\text{$\ell_{2}$}}\rangle$. Given the tree, the time complexity of computing the FT is $O((m+n)d\log(d\Phi))$ where $\Phi$ is the maximum Euclidean distance of any two feature vectors.
@@
The main computational bottleneck of FT is the construction of the tree. However, in our case, we have access to ${\mathcal{T}}$ given as prior knowledge, and thus we can skip the tree construction step. To apply tree-based EMD approximation methods, instead of building a QuadTree in FT, we first construct an augmented tree by extending the leaves of the original label tree ${\mathcal{T}}$ with samples. Specifically, since each leaf node in ${\mathcal{T}}$ corresponds to a class, we extend each leaf node by adding all samples from the corresponding class, as demonstrated in Fig.2. We set all extended edges to have edge weight $1$, and assign the instance weight $a_{i}$ for each data point. The resulting algorithm that applies Alg.2 to the augmented label tree is called Fast FlowTree (FastFT).
Key Insight
that allows FastFT to be more efficient than FT is that the tree metric between any two data samples from different classes is always identical in the augmented tree ${\mathcal{T}}$, which allows us to reduce the computation of ${\bm{P}}_{t}$ to 1d OT greedy flow matching using Alg.1. Now we can prove that Alg.1 and Alg.2 are equivalent given our constructed augmented tree in Fig.2, and the time complexity of FastFT is $O((m+n)d)$, which is optimal among all OT variants.
@@
Please refer to App.D for proof details of Theorem 3.2. Note that all OT-CPCC methods inherently need $\Omega((m+n)d)$ because we access each of $m+n$ data point once and look at each entry in $d$-dimensions. Additionally, FastFT can be applied to any label tree with no restriction on structure or edge weights. To control the importance of data points, the flexibility of FastFT can be achieved by tuning $a_{i}$.
@@
$\ell_{2}$ $O((nk+k^{2})d)$checkmark N/A
EMD $\tilde{O}(n(d+n^{2}\log n))$checkmark
Sinkhorn $\tilde{O}(n^{2}(d+I))$checkmark
TWD $\tilde{O}(n^{2}dI)$checkmark
SWD $\tilde{O}(np(d+\log n))$
FT $\tilde{O}(nd\log(d\Phi))$checkmark
FastFT $\tilde{\Theta}(nd)$checkmark
Table 1: Time complexity comparison of different CPCC methods. Let $I$ be the number of iterations for iterative methods, $p$ be the number of projections for SWD, $\Phi$ be the maximum Euclidean distance of any two feature vectors. We use $\tilde{\cdot}$ to represent a factor of $k^{2}$. For simplicity we assume $m=n$. In the batch learning setting, $d$ can be much larger than $n$ particularly for the fine-grained classification problems. See App.I for detailed analysis.
@@
To further understand the behavior of different OT approximation methods, we present experiments on synthetic datasets in Fig.3 (see App.J for more details). We first measured the computation time for different OT methods across datasets of varying sizes. Second, we evaluated the error of different OT approximation methods, along with the $\ell_{2}$ distance of the mean of two datasets, in approximating EMD. Two different scenarios were considered, either both distributions were Gaussian or non Gaussian. Under Gaussian distribution, the OT methods’ results were closely aligned with the $\ell_{2}$ method, except for SWD, which supports Prop.3.1 and implies the similarity in performance between OT and $\ell_{2}$-CPCC under Gaussian distributions. In non-Gaussian distribution scenario, notable differences emerged between OT methods and $\ell_{2}$, where SWD is closer to $\ell_{2}$ measurement. FastFT demonstrated exceptional advantage as an approximation method, outperforming Sinkhorn in both computational speed and precision.
@@ 3.3 Differentiability Analysis of OT-CPCC
Denote ${\bm{P}}^{*}$ as the optimal flow of EMD problem, then $\frac{\partial\rho_{\text{EMD}}}{\partial\theta}={\bm{P}}^{*}\cdot\frac{\partial{\bm{D}}}{\partial\theta}$.
@@
The application of Danskin’s relies on the optimality of ${\bm{P}}^{*}$. How about the other approximation methods? The differentiability is not a problem in Sinkhorn as the iterative algorithm only involves matrix multiplication. But for the SWD, TWD, FT, the approximate solutions (the flow matrix ${\bm{P}}$) for these methods are not explicitly formulated as solutions to an optimization problem. Following EMD, except for the Sinkhorn variant, our method is to treat flow matrix ${\bm{P}}$ always as a constant for these approximation methods, and stop the gradient through ${\bm{P}}$. For example, in Fast FlowTree, since the tree flow matrix ${\bm{P}}$ depends only on the structure of the tree based on external information and not on the model parameters $\theta$, we have $\frac{\partial\rho_{\text{FastFT}}}{\partial\theta}\vcentcolon={\bm{P}}_{\text{FastFT}}\cdot\frac{\partial{\bm{D}}}{\partial\theta}$.
To provide a more fine-grained understanding of the proposed method, in what follows we analyze the gradients of the proposed CPCC variants and compare them with the $\ell_{2}$-CPCC.